The blood pressure for men aged 18 to 29 is modeled with a normal density curve. The distribution has a mean of ? = 120 and a standard deviation of ? = 10. The following is the random variable distribution with the normal density curve. Use this to answer the following questions. SHOW WORK!

a. (2 point) If a man has a blood pressure of 105 how many standard deviations away from the mean is his blood pressure?
b. (2 points) What is the probability a randomly selected man will have a blood pressure of less than 105?

c. (2 points) What is the probability a randomly selected man will have a blood pressure of more than 105?
d. (2 points) Suppose a man’s blood pressure is in the 75th percentile for his age group, what is the value of the 75th percentile?
Answer #1

Ans:

a)

z=(105-120)/10

z=-1.5 which is 1.5 standard deviations below the mean.

b)

z=(105-120)/10

z=-1.5

P(z<-1.5)=normsdist(-1.5)=0.0668

c)

P(z>-1.5)=1-normsdist(-1.5)=1-0.0668=0.9332

d)

P(Z<=z)=0.75

z=normsiinv(0.75)=0.6745

P75=120+0.6745*10=126.75